# On an Erdős-Szekeres Game

### v2

@@ 3 Strategy when b=2
Suppose $a\geq b=2$. Using the representation described in Section 2, the game board can be visualized as a one-row, $(a-1)$-column grid, and each player colors the left-most unclaimed cell on their turn. Once the cell in column $(a-1)$ is claimed (by player 2 if $a$ is odd, or by player 1 if $a$ is even), the other player has no more legal moves and loses.
In terms of permutations, each player will play a new largest digit on their turn since playing any smaller digit creates a $J_{2}$ pattern and automatically loses the game. The resulting game permutation is $\pi=I_{a}$, and the game ends on the $a$th turn, leading to a loss for player 1 if $a$ is odd and a loss for player 2 if $a$ is even.
@@ 4 Strategy when b=3
We now consider the $(a,3)$-game ($a\geq 3$), and we give a winning strategy both in terms of board shading and in terms of permutation digits.
@@
The game board in this situation is a $2\times(a-1)$ grid. After each of their first $a-2$ moves, player 1 can produce a board where the first $i$ cells of row 1 are shaded and the first $i-1$ cells of row 2 are shaded, as illustrated in Figure 3. To start, player 1’s first move shades 1 cell in row 1 and 0 cells in row 2. After that, there are only two possible moves:
@@
For the end game, after $(a-2)$ moves for player 1 and $(a-3)$ moves for player 2, all cells except for $(a-1,1)$, $(a-2,2)$, and $(a-1,2)$ are shaded.
Because of Proposition 1, player 2 must shade either $(a-1,1)$ or $(a-2,2)$. In either case, player 1 can shade $(a-1,2)$ on their next move and force a player 2 loss.
In this straightforward situation, shading a cell in row 1 corresponds to playing a new left-to-right maximum, while shading a cell in row 2 corresponds to playing a non-left-to-right-maximum, in keeping with observations of maximum length $I_{a}$ and $321$-avoiders made in [6]. In terms of permutations, the winning strategy is as follows: Let $m_{1}<m_{2}<\cdots<m_{\ell}=n$ be the values of the left-to-right maxima of $\pi$ at the start of player 1’s turn.
@@
Finally, when $\pi$ has length $42a-$ 4, regardless of what player 2’s most recent move was, player 1 plays $\pi_{2a-3}=2a-4$, corresponding to cell $(a-1,2)$. On their next move, player 2 will complete either an $I_{a}$ pattern (by playing $\pi_{2a-2}\geq 2a-3$) or $J_{3}$ pattern (by playing $\pi_{2a-2}\leq 2a-4$) to lose the game.
Figure 3: A next-player-loss state for the $(a,3)$-game.
@@ 5 Strategy when b=4
We now consider the $(a,4)$-game ($a\geq 4$). This is the first case where the game board has sufficiently many rows that the same shading can be obtained by more than one permutation, and so we only describe it in terms of shaded boards. However, an interested player – thinking in terms of permutations – can consider the ordered pairs corresponding to each digit in the game so far to determine a next digit that follows the strategy given here.
@@
After this opening sequence, player 1 can always end their turn with a shaded board of one of the following two forms, illustrated in Figure 4:
@@
    Rows 1 and 2 have $k$ shaded cells and row 3 has $k-2$ shaded or eliminated cells for some $k\geq 2$.
@@
    Row 1 has $k$ shaded cells and rows 2 and 3 have $k-1$ shaded or eliminated cells for some $k\geq 2$.
@@
Note that the opening sequence produces a board that fits the first case. Now we consider each available move to player 2 and how player 1 may respond.
@@
In the first case, player 2 has four available moves: $(k+1,1)$, $(k+1,2)$, $(k-1,3)$, or $(k,3)$.
@@
  • If player 2 plays $(k+1,1)$, player 1 responds by playing $(k,3)$. This results in a board with $k+1$ shaded or eliminated cells in the first row, and $k$ shaded or eliminated cells in rows 2 and 3, which fits the second case.
  • If player 2 plays $(k+1,2)$, then $(k+1,1)$ is eliminated, and player 1 responds by playing $(k-1,3)$. This results in a board with $k+1$ shaded or eliminated cells in the first two rows and $k-1$ shaded or eliminated cells in row 3, which fits the first case.
  • If player 2 plays $(k,3)$, then $(k-1,3)$ is eliminated, and player 1 responds by playing $(k+1,1)$. This results in a board with $k+1$ shaded or eliminated cells in row 1 and $k$ shaded or eliminated cells in rows 2 and 3, which fits the second case.
  • If player 2 plays $(k-1,3)$, then player 1 responds by playing $(k+1,2)$ which eliminates $(k+1,1)$. This results in a board with $k+1$ shaded or eliminated cells in the first two rows, and $k-1$ shaded or eliminated cells in the last row, which fits the first case.
@@
Similarly, in the second case, player 2 has three available moves: $(k+1,1)$, $(k,2)$, or $(k,3)$.
  • If player 2 plays $(k+1,1)$, player 1 responds by playing $(k,3)$ which eliminates $(k,2)$. This results in a board with $k+1$ shaded or eliminated cells in row 1 and $k$ shaded or eliminated cells in rows 2 and 3, which fits the second case.
@@
  • If player 2 plays $(k,2)$, player 1 responds by playing $(k+1,2)$, which eliminates $(k+1,1)$. This results in a board with $k+1$ shaded or eliminated cells in rows 1 and 2 and $k-1$ shaded or eliminated cells in row 3, which fits the first case.
  • if player 3 plays $(k,3)$, then $(k,2)$ is eliminated, and player 1 responds by playing $(k+1,1)$. This results in a board with $k+1$ shaded or eliminated cells in row 1 and $k$ shaded or eliminated cells in rows 2 and 3, which fits the second case.
@@
The endgame begins when row 1 has $a-2$ shaded or eliminated cells.
@@
In the first case, row 2 also has $a-2$ shaded or eliminated cells and row 3 has $a-4$ shaded or eliminated cells, so the remaining unshaded region is as in Figure 4 (a). Player 2’s options are $(a-1,1)$, $(a-1,2)$, $(a-3,3)$, or $(a-2,3)$. If player 2 plays $(a-1,2)$ or $(a-2,3)$, player 1 can play $(a-1,3)$ and win the game. So player 2 will play $(a-1,1)$ or $(a-3,3)$. Whichever of these two options player 2 takes, player 1 takes the other move. This forces player 2 to play $(a-1,2)$ or $(a-2,3)$ on their next move, and player 1 takes the corner cell of $(a-1,3)$, forcing a player 2 loss.
@@
In the second case, rows 2 and 3 have $a-3$ shaded or eliminated cells, so the remaining unshaded region is as in Figure 4 (b). Player 2’s options are $(a-1,1)$, $(a-2,2)$, or $(a-2,3)$. If player 2 plays $(a-2,3)$, player 1 can play $(a-1,3)$ and win the game, or player 2 will play $(a-1,1)$ or $(a-2,2)$. In either case, player 1 plays the other move. This forces player 2 to play $(a-1,2)$ or $(a-2,3)$ on their next move, and player 1 takes the corner cell of $(a-1,3)$, forcing a player 2 loss.
(a) (b)
Figure 4: Two next-player-loss states of the $(a,4)$-game.
Similar to a $(a,3)$-game, in a $(a,4)$-game, player 1 has a clear deterministic response no matter what player 2 does. Each of these responses gets closer and closer to player 2 taking either $(a-1,2)$ or $(a-2,3)$, so that player 1 claims the bottom right corner of the board and forces a player 2 loss.
@@ 6 Strategy when b=5
Finally, we consider the $(a,5)$-game for $a\geq 5$. Although we give a winning strategy for this situation, there are more options for how player 1 selects a move leading to the final end game. The strategy given in this section was determined in an experimental matter. Since this is a 2-player combinatorial game with perfect information, when $a$ is known, every possible board shading can be labeled as a next-player win or a next-player loss via computer search. Although we are limited to small values of $a$ for a complete computer analysis, once the computer makes this labeling of all states for several specific values of $a$, an interested human can use the computer data to conjecture a subset of positions that are next-player-loss positions regardless of $a$ and form a strategy that guarantees that regardless of player 2’s move, player 1 can respond in such a way to achieve a position in the subset. As in the previous section, there are three clear phases of the game: opening moves, midgame, and endgame.
@@
    Row 1 has an odd number of open cells and $k$ shaded or eliminated cells, while rows 2, 3, and 4 have $k-1$ shaded or eliminated cells where $k\geq 2$.
@@
    Rows 1, 2, and 3 have an even number of open cells and $k$ shaded or eliminated cells, while row 4 has $k-1$ shaded or eliminated cells where $k\geq 1$.
@@
    Rows 1 and 2 have an even number of open cells and $k$ shaded or eliminated cells, while row 3 has $k-1$ shaded or eliminated cells, and row 4 has $k-3$ shaded/eliminated cells where $k\geq 3$.
@@
    Row 1 has an even number of open cells and $k$ shaded or eliminated cells, rows 2 and 3 have $k-1$ shaded or eliminated cells, and row 4 has $k-2$ shaded/eliminated cells where $k\geq 2$.
@@
    Rows 1 and 2 have $k$ shaded or eliminated cells and non-zero open cells, while rows 3 and 4 have $k-2$ shaded or eliminated cells where $k\geq 3$.
@@
    Row 1 has $k$ shaded or eliminated cells and non-zero open cells, row 2 has $k-1$ such cells, row 3 has $k-2$ such cells, and row 4 has $k-3$ such cells where $k\geq 3$.
@@
    Row 1 has one more shaded cell than row 2 and non-zero open cells. Row 2 has at least two more shaded cells than row 3, and row 3 has one more shaded cell than row 4.
These seven states are illustrated in Figure 5 and we refer to them as the set $\widehat{S}$. A tedious computer-assisted analysis shows that if the current board is in one of the states from $\widehat{S}$ at the start of player 2’s turn, then no matter where player 2 moves, player 1 has a response that returns to a state in $\widehat{S}$.
@@
Figure 5: The seven mid-game next-player-loss states for the $(a,5)$-game
  • Both columns $a-1$ and $a-2$ have open cells. Column $a-1$ has one more open cell than column $a-2$. All other columns are completely shaded or eliminated.
  • Column $a-1$ has at least two open cells; row 4 has the same number of cells. All other cells are are shaded or eliminated.
These three states are illustrated in Figure 6 and we refer to this set of states as set $\widehat{E}$.
Figure 6: The endgame next-player-loss states for the $(a,5)$-game
We will ultimately prove Theorem 4 by a sequence of lemmas.
Lemma 1.
Proof.
Player 1’s first move is $(1,1)$. Regardless of whether player 2 chooses $(1,2)$ or $(2,1)$, player 1’s second move is $(2,2)$ which results in a $2\times 2$ shaded region.
If $a-1$ is odd, then player 1 plays whichever of $(3,2)$ and $(2,3)$ was not chosen by player 2. This results in state 3 from $\widehat{S}$.
If $a-1$ is even, then player 1 plays $(3,3)$ which results in a $3\times 3$ shaded region. Player 2 has 6 choices for what they can play in response: $(4,1)$, $(4,2)$, $(4,3)$, $(1,4)$, $(2,4)$, or $(3,4)$. If $a=5$, the board is already in a state from $\widehat{E}$ and player 1 responds in a tit-for-tat way described in Lemma 3. Otherwise, player 1 responds as described below.
If player 2 plays $(4,1)$ or $(2,4)$, player 1 should play the other of these two cells, which results in state 4 from $\widehat{S}$ on player 1’s fourth move.
If player 2 plays $(4,2)$ or $(1,4)$, player 1 should play the other of these two cells, which results in state 3 from $\widehat{S}$ on player 1’s fourth move.
If player 2 plays $(4,3)$ or $(3,4)$, player 1 should play the other of these two cells, which results in state 2 from $\widehat{S}$ on player 1’s fourth move. ∎
Next, we proceed to the midgame, by showing player 1 has a strategy to return the game to a state from $\widehat{S}$ no matter how player 2 acts. The proof we give documents the result of computer search in a manner that is, admittedly tedious, but able to be verified by an interested reader.
Lemma 2.
Proof.
In state 1, if player 2 plays in row 1, player 1 responds with the cell below to turn the board to state 5. If player 2 plays in row 2, player 1 responds in the cell to the right to to turn the board to state 5. If player 2 plays in row 3, player 1 plays in row 1 to turn the board to state 4. If player 2 plays in row 4, player 1 plays in row 3 to turn the board to state 2.
In state 2, if player 2 plays in row 1, player 1 plays in row 4 and vice versa to get to state 1. If player 2 plays in row 2, player 1 plays in row 1 to get to state 6. If player 2 plays in row 3, player 1 responds by playing in row 2 to get to state 3.
In state 3, if player 2 plays the first open cell in row 1 player 1 plays the first open cell in row 4, and vice versa to get to state 6. If player 2 plays in row 2, player 1 plays the second open cell in row 4, and vice versa to get to state 5. If player 2 plays in row 3, player 2 responds by playing the second open cell in row 4 to get to state 2.
In state 4, if player 2 plays the first open cell in row 1, player 1 responds by playing the first open cell in row 2 and vice versa to get to state 6. If player 2 plays the first open cell in row 3, player 1 responds by playing the first open cell in row 4 and vice versa to get to state 2.
In state 5, if player 2 plays in row 1, player 1 responds by playing the first open cell in row 3 to get state 6. If player 2 plays in row 2, player 1 responds by playing one cell in row 4 to return to state 5. If player 2 plays the second open cell in row 3 and there are an odd number of open cells in row 1, player 1 responds by playing in row 2 to get state 3, while if there are an even number of open cells in row 1, player 1 responds by playing in row 4 to get to state 2. If player 2 plays in the first open cell in row 3, player 1 responds by playing in row 1 to get to state 6. Finally, if player 2 plays in row 4, player 1 responds by playing in row 2 to get back to state 5.
In state 6, if player 2 plays in row 1, player 1 responds by playing the first open cell in row 2 to get to state 7. If player 2 plays in row 2, player 1 responds in row 4 to get state 5. If player 2 plays row 3 and there are an even number of open cells in row 1, player 1 plays in row 2 to get state 3. If player 2 plays row 3 and there are an odd numbers of open cells in row 1, player 1 plays the second open cell in row 4 to get state 1. If player 2 plays in row 4, player 1 plays in row 2 to get state 5.
In state 7, if player 2 plays the first open cell in row 1, player 2 plays the first open cell in row 2 and vice versa, which returns to state 7. If player 2 plays the final available open cell in row 3 and there are an odd number of open cells in row 1, player 1 responds by playing the cell below player 2’s move, resulting in state 1. On the other hand, if player 2 plays the final available open cell in row 3 and there are an even number of open cells in row 1, player 1 responds by playing the cell diagonally below player 2’s move, resulting in state 4. If player 2 plays anywhere else in row 3, player 1 responds by playing in row 4 to get to state 6 or 7. If player 2 plays one cell in row 4, player 1 responds by playing one cell in row 3 to get to state 6 or 7. ∎
@@
First notice that if we consider a move by player 2 followed by the prescribed response of player 1 in Lemma 2, at most two cells from row 1 are shaded or eliminated in that pair of turns. In fact, the only time two cells from row 1 are shaded or eliminated in the same pair of turns is starting from state 2. So, it is sufficient to follow the response prescribed in Lemma 2 until there is one open cell in row 1 for states 1, 5, 6, and 7, or until there are two open cells in row 1 for states 2, 3, and 4, and then consider how one may adapt strategies to obtain a state in $\widehat{E}$ at that point of game play.
From state 1, if there is only one open cell in row 1, this is already a state in $\widehat{E}$, so respond as in Lemma 3.
From state 2, if there are two open cells in row 1, and player 2 plays in row 1 or row 4, player 1 responds as in Lemma 2 to get to state 1 with one open cell in row 1, which is a state in $\widehat{E}$. If player 2 plays in row 2, player 1 plays the last square in row 2 to get to a state in $\widehat{E}$. If player 2 plays in row 3, player 1 responds in row 1 to get to a state in $\widehat{E}$.
@@
From state 5, if there is one open cell in row 1, and player 2 takes the last cell in row 1, then player 1 takes penultimate cell in row 3 and vice versa to get to a state in $\widehat{E}$. If player 2 takes the last cell in row 2, then player 1 takes the first square in row 3 and vice versa to get to a state in $\widehat{E}$. If player 2 moves in row 4, player 1 takes the last square in row 1 to get to a state in $\widehat{E}$.
From state 6 when there is one open cell in row 1, if player 2 moves in row 1 or row 2, player 1 takes the final cell of row 2 to get to a state in $\widehat{E}$. If player 2 moves in row 3 or row 4, player 1 takes cell $(a-3,4)$ to get to a state in $\widehat{E}$.
From state 7 when there is one open cell in row 1, if player 2 moves in row 1 or row 2, player 1 takes the final cell of row 2 to get to a state in $\widehat{E}$. If player 2 takes the cell $(a-3,3)$, then player 1 takes cell $(a-3,4)$ to get to a state in $\widehat{E}$. Otherwise if player 2 takes a different cell $(c,3)$ with $c<a-3$, then player 1 responds by taking cell $(c-1,4)$, and if player 2 plays in row 4, player 1 takes one cell in row 3, returning to state 7. ∎
@@
Once the board is in a state from $\widehat{E}$, player 1 has a tit-for-tat response to any move player 2. In particular, there are either two rows (or two columns, or one row and one column) with open cells. Without loss of generality consider the case with two rows of open cells. When player 2 takes cell $(a-1,3)$ or cell $(a-2,4)$, player 1 takes cell $(a-1,4)$ and wins the game. Until then, when player 2 takes cells from one row, player 1 takes the same number of cells from the other row. ∎
@@
The set of states $\widehat{S}$ contains states where moves are available in any of the four rows. These are convenient descriptions of families of board shadings, where the lattice path dividing the shaded and eliminated cells from the open cells is translated horizontally, depending on $k$. However, these kinds of states are only convenient descriptions of families of states for player 1 to aim for once cells have been shaded or eliminated in all but the final row, so that moves are available in all four rows, rather than restricted to the top 2 or 3 rows of the board. This means that to come up with a similar strategy for larger $b$, the starting strategy of the game will take longer, to give enough moves for the top $b-2$ rows of the board to have non-zero shading.
7 Strategy for the achievement game

### v3

@@ 3 Strategy when b=2
Suppose $a\geq b=2$. Using the representation described in Section 2, the game board can be visualized as a one-row, (a1)(a-1)-column grid, and each player colors the left-most unclaimed cell on their turn. Once the cell in column (a1)(a-1) is claimed (by player 2 if $a$ is odd, or by player 1 if $a$ is even), the other player has no more legal moves and loses.
In terms of permutations, each player will play a new largest entry on their turn since playing any smaller entry creates a $J_{2}$ pattern and automatically loses the game. The resulting game permutation is $\pi=I_{a}$, and the game ends on the $a$th turn, leading to a loss for player 1 if $a$ is odd and a loss for player 2 if $a$ is even.
@@ 4 Strategy when b=3
We now consider the $(a,3)$-game ($a\geq 3$), and we give a winning strategy both in terms of board shading and in terms of permutation entries.
@@
The game board in this situation is a $2\times(a-1)$ grid. After each of their first a2a-2 moves, player 1 can produce a board where the first $i$ cells of row 1 are shaded and the first i1i-1 cells of row 2 are shaded, as illustrated in Figure 3. To start, player 1 must shade 1 cell in row 1 and 0 cells in row 2. After that, there are only two possible moves:
@@
For the end game, after (a2)(a-2) moves for player 1 and (a3)(a-3) moves for player 2, all cells except for (a1,1)(a-1,1), (a2,2)(a-2,2), and (a1,2)(a-1,2) are shaded.
Because of Proposition 1, player 2 must shade either (a1,1)(a-1,1) or (a2,2)(a-2,2). In either case, player 1 can shade (a1,2)(a-1,2) on their next move and force a player 2 loss.
In this straightforward situation, shading a cell in row 1 corresponds to playing a new left-to-right maximum, while shading a cell in row 2 corresponds to playing a non-left-to-right-maximum, in keeping with observations of maximum length permutations avoiding $I_{a}$ and $321$, made in [6]. In terms of permutations, the winning strategy is as follows: Let $m_{1}<m_{2}<\cdots<m_{\ell}=n$ be the values of the left-to-right maxima of $\pi$ at the start of player 1’s turn.
@@
Finally, when $\pi$ has length 2a42a-4, regardless of what player 2’s most recent move was, player 1 plays $\pi_{2a-3}=2a-4$, corresponding to cell (a1,2)(a-1,2). On their next move, player 2 will complete either an $I_{a}$ pattern (by playing $\pi_{2a-2}\geq 2a-3$) or $J_{3}$ pattern (by playing $\pi_{2a-2}\leq 2a-4$) to lose the game.
Figure 3: A class $\mathcal{P}$ position for the $(a,3)$-game.
@@ 5 Strategy when b=4
We now consider the $(a,4)$-game ($a\geq 4$). This is the first case where the game board has sufficiently many rows that the same shading can be obtained by more than one permutation, and so we only describe it in terms of shaded boards. However, an interested reader – thinking in terms of permutations – can consider the ordered pairs corresponding to each entry in the game so far to determine a next entry that follows the strategy given here.
@@
After this opening sequence, player 1 can always end their turn with a board of one of the following two forms:
@@
    Rows 1 and 2 have $k$ eliminated cells and row 3 has k2k-2 eliminated cells for some $k\geq 2$.
@@
    Row 1 has $k$ eliminated cells and rows 2 and 3 have k1k-1 eliminated cells for some $k\geq 2$.
@@
Note that the opening sequence produces a board that fits the first case. Now we consider each available move for player 2 and how player 1 may respond.
@@
In the first case, player 2 has four available moves: (i) $(k+1,1)$, (ii) $(k+1,2)$, (iii) (k1,3)(k-1,3), or (iv) $(k,3)$, shown below in black.
(i) (ii) (iii) (iv)
In each of these cases, player 1 has a natural response, shown below in blue.
(i) (ii) (iii) (iv)
@@
In (ii) and (iii), player 1’s response results in a board with $k+1$ eliminated cells in the first two rows and k1k-1 eliminated cells in the third row, which matches case 1. In (i) and (iv), player 1’s response results in a board with $k+1$ eliminated cells in the first row and $k$ eliminated cells in rows 2 and 3, which matches case 2.
@@
Similarly, in the second case, player 2 has three available moves: (i) $(k+1,1)$, (ii) $(k,2)$, or (iii) $(k,3)$, shown below in black.
(i) (ii) (iii)
In each of these cases, player 1 again has a natural response, shown below in blue.
(i) (ii) (iii)
@@
In (ii), player 1’s response results in a board with $k+1$ eliminated cells in the first two rows and k1k-1 eliminated cells in the third row, which matches case 1. In (i) and (iii), player 1’s response results in a board with $k+1$ eliminated cells in the first row and $k$ eliminated cells in rows 2 and 3, which matches case 2.
@@
The endgame begins at the start of player 2’s turn when row 1 has a2a-2 eliminated cells, shown in two cases below.
(i) (ii)
@@
In case (i), player 2’s options are (a1,1)(a-1,1), (a1,2)(a-1,2), (a3,3)(a-3,3), or (a2,3)(a-2,3). If player 2 plays (a1,2)(a-1,2) or (a2,3)(a-2,3), player 1 can play (a1,3)(a-1,3) and win the game. So player 2 will play (a1,1)(a-1,1) or (a3,3)(a-3,3). Whichever of these two options player 2 takes, player 1 takes the other move. This forces player 2 to play (a1,2)(a-1,2) or (a2,3)(a-2,3) on their next move, and player 1 takes the corner cell of (a1,3)(a-1,3), forcing a player 2 loss.
@@
In case (ii) player 2’s options are (a1,1)(a-1,1), (a2,2)(a-2,2), or (a2,3)(a-2,3). If player 2 plays (a2,3)(a-2,3), player 1 can play (a1,3)(a-1,3) and win the game, or player 2 will play (a1,1)(a-1,1) or (a2,2)(a-2,2). In either case, player 1 plays the other move. This forces player 2 to play (a1,2)(a-1,2) or (a2,3)(a-2,3) on their next move, and player 1 takes the corner cell of (a1,3)(a-1,3), forcing a player 2 loss.
@@ 6 Strategy when b=5
Finally, we consider the $(a,5)$-game for $a\geq 5$. Although we give a winning strategy for this situation, there are more options for how player 1 selects a move leading to the final endgame. The strategy given in this section was determined in an experimental manner. Since this is a two-player combinatorial game with perfect information, when $a$ is known, every possible game state can be labeled as a next-player win (in class $\mathcal{N}$) or a next-player loss (in class $\mathcal{P}$) via computer search. Although we are limited to small values of $a$ for a complete computer analysis, once the computer makes this labeling of all states for several specific values of $a$, an interested human can use the computer data to conjecture a subset of positions that are in class $\mathcal{P}$ regardless of $a$ and form a strategy that guarantees that regardless of player 2’s move, player 1 can respond in such a way to achieve a position in the subset. As in the previous section, there are three clear phases of the game: opening moves, midgame, and endgame.
@@
    Row 1 has an odd number of open cells and $k$ eliminated cells, while rows 2, 3, and 4 have k1k-1 eliminated cells where $k\geq 2$.
    odd
@@
    Rows 1, 2, and 3 have an even number of open cells and $k$ eliminated cells, while row 4 has k1k-1 eliminated cells where $k\geq 1$.
    even
@@
    Rows 1 and 2 have an even number of open cells and $k$ eliminated cells, while row 3 has k1k-1 eliminated cells, and row 4 has k3k-3 eliminated cells where $k\geq 3$.
    even
@@
    Row 1 has an even number of open cells and $k$ eliminated cells, rows 2 and 3 have k1k-1 eliminated cells, and row 4 has k2k-2 eliminated cells where $k\geq 2$.
    even
@@
    Rows 1 and 2 have $k$ eliminated cells and non-zero open cells, while rows 3 and 4 have k2k-2 eliminated cells where $k\geq 3$.
@@
    Row 1 has $k$ eliminated cells and non-zero open cells, row 2 has k1k-1 such cells, row 3 has k2k-2 such cells, and row 4 has k3k-3 such cells where $k\geq 3$.
@@
    Row 1 has one more eliminated cell than row 2 and non-zero open cells. Row 2 has at least two more eliminated cells than row 3, and row 3 has one more eliminated cell than row 4.
    $\cdots$
We refer to these seven states as the set $\widehat{S}$. A tedious computer-assisted analysis shows that if the current board is in one of the states from $\widehat{S}$ at the start of player 2’s turn, then no matter where player 2 moves, player 1 has a response that returns to a state in $\widehat{S}$.
  1. 1.
    Both columns a1a-1 and a2a-2 have open cells. Column a1a-1 has one more open cell than column a2a-2. All other columns are completely eliminated.
  2. 2.
    Column a1a-1 has at least two open cells; row 4 has the same number of open cells. All other cells are are eliminated.
  3. 3.
We refer to this set of three states as set $\widehat{E}$.
We will ultimately prove Theorem 4 by a sequence of lemmas.
Lemma 1.
Proof.
Player 1’s first move is $(1,1)$. Regardless of whether player 2 chooses $(1,2)$ or $(2,1)$, player 1’s second move is $(2,2)$ which results in a $2\times 2$ eliminated region.
If a1a-1 is odd, then player 1 plays whichever of $(3,2)$ and $(2,3)$ was not chosen by player 2. This results in State 3 from $\widehat{S}$.
If a1a-1 is even, then player 1 plays $(3,3)$ which results in a $3\times 3$ eliminated region. Player 2 has six choices for what they can play in response: $(4,1)$, $(4,2)$, $(4,3)$, $(1,4)$, $(2,4)$, or $(3,4)$. If $a=5$, the board is already in a state from $\widehat{E}$ and player 1 responds in a tit-for-tat way described in Lemma 3. Otherwise, player 1 responds as described below.
If player 2 plays $(4,1)$ or $(2,4)$, player 1 plays the other of these two cells, which results in State 4 from $\widehat{S}$ on player 1’s fourth move.
If player 2 plays $(4,2)$ or $(1,4)$, player 1 plays the other of these two cells, which results in State 3 from $\widehat{S}$ on player 1’s fourth move.
If player 2 plays $(4,3)$ or $(3,4)$, player 1 plays the other of these two cells, which results in State 2 from $\widehat{S}$ on player 1’s fourth move. ∎
Next, we proceed to the midgame, by showing player 1 has a strategy to return the game to a state from $\widehat{S}$ no matter how player 2 acts. The proof we give documents the result of computer search in a manner that is admittedly tedious, but able to be verified by an interested reader.
Lemma 2.
Proof.
In State 1, player 2 has four possible moves, shown below, followed by player 1’s responses.
even
odd
odd
odd
even
even
State 5 State 5 State 4 State 2
In State 2, player 2 has four possible moves, shown below, followed by player 1’s responses.
odd
odd
odd
even
odd
even
odd
State 1 State 6 State 3 State 1
In State 3, player 2 has five possible moves, shown below, followed by player 1’s responses.
odd
@@
even
State 6 State 5 State 2 State 5 State 6
In State 4, player 2 has four possible moves, shown below, followed by player 1’s responses.
odd
even
even
even
even
even
State 6 State 6 State 2 State 2
In State 5, player 2 has five possible moves, shown below, followed by player 1’s responses. In one of these cases, player 1’s response depends on the parity of the number of open cells in row 1.
odd
even
even
even
State 6 State 5 State 3 State 2 State 6 State 5
In State 6, player 2 has four possible moves, shown below, followed by player 1’s responses. In one of these cases, player 1’s response depends on the parity of the number of open cells in row 1.
even
odd
even
odd
State 7 State 5 State 3 State 1 State 5
Finally, recall that in State 7, row 1 has one more eliminated cell than row 2, and row 3 has one more eliminated cell than row 4, but row 3 has at least two more eliminated cells than row 3, as illustrated below.
@@
In general, if player 2 plays in row 1, player 1 responds in row 2 and vice versa to remain in State 7. Similarly, if player 2 plays in row 4, player 1 responds with one cell in row 3 to get to State 6 or State 7. The interesting situation is when player 2 plays in row 3, since there is more than one open cell in row 3 to choose from. If player 2 plays anywhere but rightmost possible cell of row 3, player 1 plays in row 4 in such a way as to return to State 6 or 7. The remaining case is when player 2 plays the final possible cell in row 3, which we consider in two cases below, depending on the parity of open cells in row 1.
[…]
State 1 State 4
@@
First notice that if we consider a move by player 2 followed by the prescribed response of player 1 in Lemma 2, at most two cells from row 1 are eliminated in that pair of turns. In fact, the only time two cells from row 1 are eliminated in the same pair of turns is starting from State 2. So, it is sufficient to follow the response prescribed in Lemma 2 until there is one open cell in row 1 for States 1, 5, 6, and 7, or until there are two open cells in row 1 for States 2, 3, and 4, and then consider how one may adapt strategies to obtain a state in $\widehat{E}$ at that point of game play.
In State 1, if there is only one open cell in row 1, this is already a state in $\widehat{E}$.
From State 2, if there are two open cells in row 1, and player 2 plays in row 1 or row 4, player 1 responds as in Lemma 2 to get to State 1 with one open cell in row 1, which is a state in $\widehat{E}$. If player 2 plays in row 2, player 1 plays the last cell in row 2 to get to a state in $\widehat{E}$. If player 2 plays in row 3, player 1 responds in row 1 to get to a state in $\widehat{E}$.
@@
From State 5, if there is one open cell in row 1, and player 2 takes the last cell in row 1, then player 1 takes penultimate cell in row 3 and vice versa to get to a state in $\widehat{E}$. If player 2 takes the last cell in row 2, then player 1 takes the first cell in row 3 and vice versa to get to a state in $\widehat{E}$. If player 2 moves in row 4, player 1 takes the last cell in row 1 to get to a state in $\widehat{E}$.
From State 6 when there is one open cell in row 1, if player 2 moves in row 1 or row 2, player 1 takes the final cell of row 2 to get to a state in $\widehat{E}$. If player 2 moves in row 3 or row 4, player 1 takes cell (a3,4)(a-3,4) to get to a state in $\widehat{E}$.
From State 7 when there is one open cell in row 1, if player 2 moves in row 1 or row 2, player 1 takes the final cell of row 2 to get to a state in $\widehat{E}$. If player 2 takes the cell (a3,3)(a-3,3), then player 1 takes cell (a3,4)(a-3,4) to get to a state in $\widehat{E}$. Otherwise if player 2 takes a different cell $(c,3)$ with c<a3c<a-3, then player 1 responds by taking cell (c1,4)(c-1,4), and if player 2 plays in row 4, player 1 takes one cell in row 3, returning to State 7. ∎
@@
Once the board is in a state from $\widehat{E}$, player 1 has a tit-for-tat response to any move from player 2. In particular, there are either two rows (or two columns, or one row and one column) with open cells. Without loss of generality consider the case with two rows of open cells. When player 2 takes cell (a1,3)(a-1,3) or cell (a2,4)(a-2,4), player 1 takes cell (a1,4)(a-1,4) and wins the game. Until then, when player 2 takes cells from one row, player 1 takes the same number of cells from the other row. ∎
@@
The set of states $\widehat{S}$ contains states where moves are available in any of the four rows. These are convenient descriptions of families of board shadings, where the lattice path dividing the eliminated cells from the open cells is translated horizontally, depending on $k$. However, these kinds of states are only convenient descriptions of families of states for player 1 to aim for once cells have been eliminated in all but the final row, so that moves are available in all four rows, rather than restricted to the top 2 or 3 rows of the board. This means that to come up with a similar strategy for larger $b$, the starting strategy of the game will take longer, to give enough moves for each of the top b2b-2 rows of the board to have non-zero eliminated cells.
7 Other Directions
We conclude with comments on several other directions of interest.
7.1 Strategy for normal play